how do i rewrite the following using a single exponent?
4^(10m-6)/4^(9m-7)=4

Answer by lwsshak3(11628)

how do i rewrite the following using a single exponent?
4^(10m-6)/4^(9m-7)=4
4^(10m-6)*4^-(9m-7)=4
4^(10m-6)-(9m-7)=4
4^(10m-6-9m+7=4
4^(m+1)=4